### v1

@@ 2.1 Softmax, Safe Softmax, Online Softmax
The softmax function takes as input a vector $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})\in\mathbb{R}^{N}$ and returns a normalized vector $\sigma(\mathbf{x})\in\mathbb{R}^{N}$ over the NN components. Formally, the $i$-th component of the softmax function $\sigma(\mathbf{x})_{i}$ is defined as:
@@
When computing the denominator of Eq. 2, a naive implementation uses two serial loops to compute the max and sum. This implementation is inefficient for large NN because it prevents efficient tiling of the computation.
@@ 2.2 Attention, FlashAttention, FlexAttention
Softmax is applied row-wise to the attention matrix $\frac{QK^{\top}}{\sqrt{d_{k}}}$, producing normalized attention weights that sum to 11 over the key dimension for each query. Listing 1 shows the PyTorch implementation of scaled dot-product attention (with attn_mask=None). We refer to this as Vanilla Attention. Computing the attention matrix or attn_scores requires $\mathcal{O}(n^{2})$ time and memory, which becomes a bottleneck for long sequences.
@@
7 # Set INF when mask is True
@@
12 # Set INF when mask is False
@@ 3.3 Algebraic Transformation of Reductions
Let AA be a set with a binary operation $\oplus$, and let $B$ be a set with a binary operation $\otimes$. A function $f:A{\rightarrow}B$ is said to be a homomorphism if for all $a_{1},a_{2}\in A$, $f(a_{1}\oplus a_{2})=f(a_{1})\otimes f(a_{2})$.
In the context of softmax, $A=B=\mathbb{R}$ (the set of real numbers), and $\oplus$ and $\otimes$ are addition (++) and multiplication ($\times$) of real numbers. The function $f(x)=e^{x}$ is a homomorphism because $f^{a+b}=f^{a}\times f^{b}$. To generalize the online softmax construction, we need the set AA with operations $\oplus$ and $\otimes$ to satisfy the axioms of a ring:
@@
In the stable softmax algorithm (Alg. 1), let us denote the sequence of $m$ values produced by the first loop by $m[1..N]$ and let $m[0]=0$ by definition. The sequence of $d$ values produced by the second loop, which we denote by dsds, is expressed abstractly by the following recurrence in which the elements of dsds are members of a ring AA and $E{:}A{\rightarrow}A$ is a homomorphism.
@@
It is easy to show by induction that dsds can be expressed in closed-form by the following expression in which $\bigoplus$ stands for the application of the associative operation $\oplus$ to a set of elements of AA.
@@
The online softmax algorithm (Alg. 2) computes a different sequence, denoted by dodo, that can be expressed abstractly as shown below.
@@
The dsds and dodo sequences will be different in general, but we show that dodo can be expressed in closed-form by the following expression.
@@ 3.5 Structural Fusion with Tiling-Aware Dimension Elimination
These kernels are incompatible under standard fusion rules. Without tiling, these sketches are also incompatible for the producer-consumer fusion described previously; the producer’s p-dimension NN does not match the consumer’s p-dimension $P$. Fusing them would require materializing the entire intermediate tensor $C$ into global memory.
@@
using our generalized fusion rule with dimension demotion. The dimension NN from the producer’s output is consumed directly as an r-dimension by the consumer on-the-fly within a single fused kernel, completely avoiding the materialization of the intermediate tensor $C$. This tiling-aware transformation allows FlashLight to fuse complex, multi-stage computations that have historically remained out of reach for the PyTorch compiler.

### v2

@@ 2.1 Softmax, Safe Softmax, Online Softmax
The softmax function takes as input a vector $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})\in\mathbb{R}^{N}$ and returns a normalized vector $\sigma(\mathbf{x})\in\mathbb{R}^{N}$ over the $N$ components. Formally, the $i$-th component of the softmax function $\sigma(\mathbf{x})_{i}$ is defined as:
@@
When computing the denominator of Eq. 2, a naive implementation uses two serial loops to compute the max and sum. This implementation is inefficient for large $N$ because it prevents efficient tiling of the computation.
@@ 2.2 Attention, FlashAttention, FlexAttention
Softmax is applied row-wise to the attention matrix $\frac{QK^{\top}}{\sqrt{d_{k}}}$, producing normalized attention weights that sum to $1$ over the key dimension for each query. Listing 1 shows the PyTorch implementation of scaled dot-product attention (with attn_mask=None). We refer to this as Vanilla Attention. Computing the attention matrix or attn_scores requires $\mathcal{O}(n^{2})$ time and memory, which becomes a bottleneck for long sequences.
@@
7 # Set -INF when mask is True
@@
12 # Set -INF when mask is False
@@ 3.3 Algebraic Transformation of Reductions
Let $A$ be a set with a binary operation $\oplus$, and let $B$ be a set with a binary operation $\otimes$. A function $f:A{\rightarrow}B$ is said to be a homomorphism if for all $a_{1},a_{2}\in A$, $f(a_{1}\oplus a_{2})=f(a_{1})\otimes f(a_{2})$.
In the context of softmax, $A=B=\mathbb{R}$ (the set of real numbers), and $\oplus$ and $\otimes$ are addition (++) and multiplication ($\times$) of real numbers. The function $f(x)=e^{x}$ is a homomorphism because $f^{a+b}=f^{a}\times f^{b}$. To generalize the online softmax construction, we need the set $A$ with operations $\oplus$ and $\otimes$ to satisfy the axioms of a ring:
@@
In the stable softmax algorithm (Alg. 1), let us denote the sequence of $m$ values produced by the first loop by $m[1..N]$ and let $m[0]=0$ by definition. The sequence of $d$ values produced by the second loop, which we denote by $ds$, is expressed abstractly by the following recurrence in which the elements of $ds$ are members of a ring $A$ and $E{:}A{\rightarrow}A$ is a homomorphism.
@@
It is easy to show by induction that $ds$ can be expressed in closed-form by the following expression in which $\bigoplus$ stands for the application of the associative operation $\oplus$ to a set of elements of $A$.
@@
The online softmax algorithm (Alg. 2) computes a different sequence, denoted by $do$, that can be expressed abstractly as shown below.
@@
The $ds$ and $do$ sequences will be different in general, but we show that $do$ can be expressed in closed-form by the following expression.
@@ 3.5 Structural Fusion with Tiling-Aware Dimension Elimination
These kernels are incompatible under standard fusion rules. Without tiling, these sketches are also incompatible for the producer-consumer fusion described previously; the producer’s p-dimension $N$ does not match the consumer’s p-dimension $P$. Fusing them would require materializing the entire intermediate tensor $C$ into global memory.
@@
using our generalized fusion rule with dimension demotion. The dimension $N$ from the producer’s output is consumed directly as an r-dimension by the consumer on-the-fly within a single fused kernel, completely avoiding the materialization of the intermediate tensor $C$. This tiling-aware transformation allows FlashLight to fuse complex, multi-stage computations that have historically remained out of reach for the PyTorch compiler.